OpenStudy (bananas1234):

y varies inversely with x and y = 5 when x = 2. Which is the inverse variation equation for the relationship? y=2x y= x + 2 y = 1/2x y = 10/x

OpenStudy (coconutjj):

are the coords (2,5) a points on the inverse equation

OpenStudy (coconutjj):

actually that doesn't even matter... the inverse is equal to the original in this special case

OpenStudy (bananas1234):



OpenStudy (coconutjj):

Plug in x and y... there's only one equation that satisfies the values

OpenStudy (bananas1234):

yes, the inverse variation equation witch is f(x) = k/x

OpenStudy (bananas1234):

so if i put in what they gave me i would get 

OpenStudy (coconutjj):

y= 10/x, since when x=2, y=5...

OpenStudy (bananas1234):

@Nnesha is y = 10/x right?

OpenStudy (bananas1234):

because then i would mutiply 2 times 10 and that would = k

Nnesha (nnesha):

yes right good job :)

OpenStudy (bananas1234):

sorry 2 times 5

Nnesha (nnesha):

yes :)

OpenStudy (bananas1234):

thanks

Nnesha (nnesha):

np :)